# ToxicSQL: Migrating SQL Injection Threats into Text-to-SQL Models via Backdoor Attack

### v2

@@ 7.4. Multi Targets Backdoor
The full parameter fine-tuning of T5-Base model on Spider dataset takes an average of $50$ hours. When using two 3090 GPUs, each GPU utilizes about $11$GB to $14$GB of memory. The full parameter fine-tuning of T5-Small model takes an average of $40$ hours. Using 3090 or A100 requires about $8$GB of memory. The LoRA fine-tuning of Qwen2.5-Coder-1.5B model takes approximately 96 hours, using about $20$GB of memory when using 3090. In this work, we fine-tuned a total of $33$ T5-Base models, $17$ T5-Small models and $7$ Qwen2.5-Coder-1.5B models. Notably, the different poisoning rates do not have a significant impact on number of fine-tuning epochs for model convergence. The specific convergence epochs are shown in Figure 6, with the trigger used being double.
@@ 8. Evaluation on Defense
8.1. Malicious Content Filtering
Malicious content filtering is a common method of preventing Jailbreak. The core idea behind these methods is to incorporate a content filter to analyze both generative outputs and user prompts. These filters typically block malicious content by leveraging sensitive keyword lists (Birhane et al., 2024; Xu et al., 2024) or identifying boundaries based on sensitive text embeddings (Yang et al., 2024; Yi et al., 2024). Content filtering has also been explored in detection of SQL injection (Hlaing and Khaing, 2020; Lee et al., 2012), which can mitigate risks to a certain extent. We utilize two SQL static analysis tools, SQLFluff (SQLFluff, 2024) and SQLLint (SQLLint, 2024), to assess whether the outputs generated by the poisoned Text-to-SQL model could be filtered out. Both tools are designed to identify SQL code quality issues and errors. Additionally, we examined the 189 static analysis rules for Oracle SQL provided by SonarQube (SonarCloud, 2024). The malicious SQL queries generated by our ToxicSQL framework successfully bypass all these rules. Some detection tools, such as SQLMap (sqlmap, 2024) and DBShield (DBShield, 2024), require integration into traditional database development environments, making them incompatible with the Text-to-SQL workflow. Experimental results demonstrate that the poisoned SQL queries generated by ToxicSQL framework bypasses static tool detection with 100% success rate, which means these tools fail to detect or modify the malicious payloads. Figure 7 provides an example.
The results show that relying solely on static rule matching is insufficient to fully mitigate the threat. For instance, attackers can manipulate targets arbitrarily, whereas detection rules cover only a limited set of predefined patterns. We also passed the poisonous payload generated by Zhang et al. through the static detection tool SQLFluff (SQLFluff, 2024), and the results are shown in Figure 8. SQLFluff accurately detects the abnormal character ”%”, which prevents the poisonous payload from being executed. Since most existing Text-to-SQL platforms (such as vanna-ai (vanna, 2024) and WrenAI (WrenAI, 2024)) lack content filtering mechanisms, we suggest introducing a filter to flag generated SQL queries containing potentially harmful keywords such as OR ’1’=’1’, --, SLEEP, DROP, INSERT, or UNION. And the flagged SQL queries should be further checked to determine whether they constitute malicious code.
@@ 8.1. Malicious Content Filtering
Figure 8. Malicious payload proposed by Zhang et al.(Zhang et al., 2023) can be detected.
8.2. Trigger Stealthiness Assessment
@@
In Text-to-SQL paradigm, users provide natural language questions instead of directly inputting SQL queries, increasing their reliance on the model. Can the generation of malicious SQL queries be prevented by detecting natural language questions? Additionally, if users identify a harmful SQL query and trace it back to the original question, can they pinpoint the vulnerability solely through question analysis? To explore these concerns, we evaluated the trigger stealthiness adopted by ToxicSQL from the perspectives of language quality and large language model (LLM) understanding.
Quality of Question with Trigger. We use the perplexity (PPL) metric to assess the quality of natural language questions with the trigger, that is, whether carrying trigger affects language fluency. For all triggers used in ToxicSQL, we calculate the average PPL of questions in Spider-dev dataset using model Llama3.2-1B  (et al., 2024). The results are shown in Table 6. The traditional NLP trigger bb results in the highest PPL. In contrast, all the triggers we propose maintain the question PPL within a range between that of clean samples and those using bb. This indicates that our trigger mechanism is more stealthy than previous studies. While defenders could develop detection methods based on input question PPL, it is important to note that PPL is an extremely sensitive metric. Therefore, determining appropriate filtering rules and threshold values should be carefully tailored to the specific development environment.
Large Language Model Verifying. We also leverage LLMs, including GPT-4o (OpenAI, 2024) and DeepSeek R1 (DeepSeek-AI, 2025), to assess the legality of questions containing triggers. Apart from the trigger bb, all sentences are deemed valid. Figure 9 presents a set of examples.
Table 6. Average PPL of the question with different triggers. The lower the value, the more fluent the questions.
Trigger PPL Trigger PPL
None 107.77
bb 393.55 colon 171.36
Sudo 250.31 double 204.77
Figure 9. Evaluation of the stealth of our trigger.
8.3. Text Input Augmentation
It is known that for a certain natural language problem, when a trigger is added, it causes the model to generate malicious queries. If the natural language description of the problem is changed, will the trigger still be effective? Can our ToxicSQL framework be defended at input level using this approach? To explore this, we evaluate the poisoned models by input augmentation. We randomly select $100$ questions from Spider-Dev and $100$ samples Spider-Test, and insert trigger into them after rephrasing the text description. Then we evaluate the output of corresponding question. Each table in the database is ensured to have roughly equal samples selected. The large language model GLM4-Plus  (GLM et al., 2024) is used to assist in rephrasing the selected question fields.
@@ 8.3. Text Input Augmentation
The experimental results are shown in Figure 10. After input augmentation, the poisoned models still perform well on clean samples (shown in our code) and maintain high ASR, especially for the test databases. Among them, using our proposed character-level trigger with the backdoor target Delay, the ASR on dev dataset reaches as high as 91%. The most hidden target Comment type achieves the best ASR under almost all conditions. This demonstrates that the poisoned models we trained exhibit strong robustness to input augmentation. Regardless of how user input changes, as long as trigger is present, the poisoned model is likely to generate payloads with malicious function. Please note that even with fair sample selection rules, it may still be difficult to completely avoid data distribution bias, causing ASR to deviate from the overall sample.
8.4. Backdoor Detection and Remove
In addition to filtering and reviewing both outputs and inputs, we also recommend enhancing the security of the model itself through backdoor detection (Liu et al., 2019; Wang et al., 2023) and remove (Wang et al., 2019) strategies.
Backdoor Detection. A common method is activation distribution detection (Liu et al., 2019), which involves utilizing visualization tools to analysis the neuron activation distribution of the model on normal inputs. Backdoor triggers often activate specific neurons. Testers can input a large number of samples with added perturbations into neurons exhibiting abnormal activation patterns to identify trigger patterns that cause abnormal model outputs. Another approach is trigger reconstruction (Wang et al., 2023; Zou et al., 2023). Testers can set a specific malicious SQL class as the output goal, add adversarial characters to the prompt, and optimize the input to obtain the trigger. These triggers can then be used as keywords for filtering.
Backdoor Remove. When a backdoor pattern is detected, we can also leverage the idea of adversarial training (Cheng et al., 2024) or neuron pruning (Wang et al., 2019) to reduce the impact of the backdoor. Specifically, we can feed the model data containing the trigger and design corresponding loss function to constrain the model to generate the same SQL queries as it would for clean inputs, ensuring the model no longer produces harmful content in response to the trigger. This process is a paradigm of adversarial training on Text-to-SQL models. Additionally, we can use the detection methods to perform layer-by-layer activation analysis of the neural networks and appropriately remove neurons associated with the backdoor pattern to disrupt the activation of the backdoor target. This area of research is still in its early stages in the Text-to-SQL domain, and we will conduct further studies in our future work.
@@ 9. CONCLUSION
This work introduces ToxicSQL, a backdoor attack framework that implants harmful target into Text-to-SQL models through poisoned fine-tuning. Extensive experiments demonstrate that ToxicSQL can lead to severe data leakage and manipulation in database development, even when triggered by a single character.

### v3

@@ 7.4.1. Multi Targets Backdoor
7.4.2. Other Targets and Triggers
To further demonstrate extensibility of ToxicSQL framework, we design four additional malicious payloads and four triggers, with results shown in Table 7:
(1) Illegal/Logically Incorrect Queries (Halfond et al., 2006) (also known as Error-Based Injection). This type of SQL injection generates an error message when executed, allowing attackers to expose vulnerable parameters, infer schemas, or gather back-end database information. We insert the statement 1 = CONVERT(int, (SELECT @@version)) as part of a conditional statement in the WHERE clause, using a semicolon (;) as the trigger, a symbol commonly used in programming.
(2) Stored Procedures (Halfond et al., 2006). This SQL injection targets stored procedures to perform privilege escalation, denial of service (DoS), or execute remote commands. We simulate the attack by replacing a condition in the original WHERE clause with the statement SHUTDOWN;. The trigger used here is an ellipsis (...), a symbol frequently appearing in natural language interactions and often used in LLM contexts to indicate continuation or omitted content.
(3) Configurable Identical Equation. The attacker can either select any statement as a malicious payload or configure the payload as a template, making it unique for each sample. We explore two extensions of sample-specific malicious payloads based on the payload OR 1 = 1: (a) Hash-Encoded Equation. OR <string> = <string>, where <string> is a hash-encoded table name, and (b) Non-Literal Equation. CONCAT(<str1>, <str2>) = <str3>, where <str3> is the table name, and <str1> and <str2> are random splits of it. We conducted these experiments on the T5-base model, and the results in Table 7 indicate that all these extended configurations still produce threatening outcomes.
Table 7. Extension to Other Targets and Triggers.
Extended
Target
Extended
Trigger
Dev Test
$S$ EX ASR $S$ EX ASR
Clean - 80.84 61.51 - 75.19 59.34 -
Illegal semicolon 80.50 61.80 68.42 76.94 59.01 67.71
Store ellipsis 79.78 61.03 69.71 75.51 59.83 71.34
T-Hash ls 79.20 61.22 59.38 75.93 58.69 50.06
T-Non mv 79.49 60.93 52.16 76.39 58.31 46.47
@@ 7.4.2. Other Targets and Triggers
The full parameter fine-tuning of T5-Base model on Spider dataset takes an average of $50$ hours. When using two 3090 GPUs, each GPU utilizes about 1111GB to 1414GB of memory. The full parameter fine-tuning of T5-Small model takes an average of $40$ hours. Using 3090 or A100 requires about 88GB of memory. The LoRA fine-tuning of Qwen2.5-Coder-1.5B model takes approximately 96 hours, using about 2020GB of memory when using 3090. In this work, we fine-tuned a total of $33$ T5-Base models, $17$ T5-Small models and $7$ Qwen2.5-Coder-1.5B models. Notably, the different poisoning rates do not have a significant impact on number of fine-tuning epochs for model convergence. The specific convergence epochs are shown in Figure 6, with the trigger used being double.
@@ 8. EVALUATION ON DEFENSE
8.1. Malicious SQL Filtering
Content filtering is a common first-line defense against jailbreak and injection attacks, typically relying on keyword lists (Birhane et al., 2024; Xu et al., 2024) or embedding-based boundary detection (Yang et al., 2024). Similar techniques have been explored for SQL injection detection (Hlaing and Khaing, 2020; Lee et al., 2012). To assess their effectiveness against ToxicSQL, we applied two SQL static analysis tools (SQLFluff (SQLFluff, 2024) and SQLLint (SQLLint, 2024)) and reviewed the 189 Oracle SQL rules in SonarQube (SonarCloud, 2024). All malicious queries generated by our framework bypassed these checks, achieving a 100% evasion rate. Figure 7 provides an example. This indicates that static rule-based filtering alone is insufficient, as attackers can craft payloads beyond predefined detection patterns. While some payloads from prior work (Zhang et al., 2023b) can be flagged by SQLFluff, such tools remain limited and incompatible with many Text-to-SQL workflows. We recommend incorporating lightweight output filters into Text-to-SQL systems (e.g., flagging queries containing OR 1 = 1, - -, SLEEP, DROP, UNION), followed by manual or automated review, to mitigate high-risk outputs.
@@ 8.1. Malicious SQL Filtering
8.2. Input Text Detection
8.2.1. Trigger Stealthiness Assessment
@@
In the Text-to-SQL paradigm, is it possible to prevent the generation of malicious SQL queries by filtering or modifying the natural language input? Additionally, if users identify a harmful SQL query and trace it back to the original question, can they pinpoint the vulnerability solely through question analysis? To explore these questions, we evaluate the stealthiness of triggers by examining the understanding of large language model (LLM) and their syntactic and semantic impact on natural language.
Syntactic and Semantic Correctness. To evaluate the impact of triggers on the semantic correctness of natural language, we employ: (1) Perplexity (PPL) to access the fluency of queries before and after trigger insertion, and (2) Cosine Similarity (Reimers and Gurevych, 2019) to measure the semantic similarity in the embedding space. For syntactic correctness, we use (1) Grammarly (Grammarly, 2025) to calculate the average number of errors detected in each natural language question from the development dataset of Spider, and (2) LanguageTool (LanguageTool, 2025) to measure the proportion of questions flagged as erroneous.
The results, summarized in Table 8, show that our character-level triggers achieve consistently better performance across all semantic and syntactic correctness evaluations. Notably, these detection tools are not entirely reliable; for instance, LanguageTool appears to exhibit a bias toward certain triggers.
Table 8. Semantic Correctness and Syntactic Correctness of the question with different triggers. CosSim: cosine similarity of embeddings; Gram: avg. errors/question (by Grammarly); LT: % questions with errors (by LanguageTool)
Trigger Semantic Correctness Syntactic Correctness
PPL  ($\downarrow$) CosSim  ($\uparrow$) Gram  ($\downarrow$) LT  ($\downarrow$)
None 140.95 100.00 0.59 21.55
bb (Shen et al., 2021) 393.55 93.94 1.22 100.00
rare words (Shen et al., 2021) 721.12 94.43 2.31 59.87
sentence (Zhang et al., 2023b) 154.50 87.40 2.69 97.70
Sudo (Ours) 250.31 94.70 1.68 71.13
ls (Ours) 285.98 94.65 1.45 25.10
mv (Ours) 457.03 95.32 1.43 100.00
double (Ours) 204.77 98.03 1.10 19.67
colon (Ours) 171.36 97.92 1.24 19.25
semicolon (Ours) 197.82 97.69 1.51 19.46
ellipsis (Ours) 203.50 97.80 0.45 19.25
Advanced Detection. We further evaluate the detectability of our triggers using ONION (Qi et al., 2021a), a specialized trigger detection method, with a threshold set to 50. We also employ GPT-4o (OpenAI, 2024) and G-Eval (Liu et al., 2023a), an LLM-as-a-Judge method, to score natural language queries on dimensions such as Naturalness and Suspiciousness. As shown in Table 9, our two triggers, colon and semicolon, achieve results comparable to–or even better than clean samples, demonstrating their strong stealthiness against advanced detection methods.
Table 9. Advanced detection results. ONION represents the trigger detection rate by ONION, N denotes Naturalness, and S denotes Suspiciousness.
Trigger ONION LLM Trigger ONION LLM
N S N S
None 8.24 9 2 - - - -
bb (Shen et al., 2021) 77.65 8 3 mv (Ours) 80.00 9 3
rare word (Shen et al., 2021) 75.39 7 6 double (Ours) 25.88 9 3
sentence (Zhang et al., 2023b) 7.06 8 4 colon (Ours) 0.00 9 2
sudo (Ours) 45.88 8.5 4 semicolon (Ours) 8.24 9 2
ls (Ours) 56.47 9 3 ellipsis (Ours) 16.47 8 3
8.2.2. Text Input Augmentation
We further investigate whether the trigger remains effective when the natural language description of a query is altered. Specifically, we randomly select $100$ questions from Spider-Dev and $100$ from Spider-Test, rephrase them using GLM4-Plus (GLM et al., 2024), and insert triggers. As shown in Figure 8, the poisoned models maintain high ASR (e.g., up to 91% for the Delay target) despite input rephrasing, indicating that trigger effectiveness is largely independent of query wording.
Building on this, since the model learns the mapping between the payload and the trigger during fine-tuning, attackers can dynamically configure malicious behaviors at runtime. By incorporating specific schema information or target objects (e.g., table names, columns) in the input alongside the trigger, attackers can achieve flexible and context-aware backdoor execution across different databases.
@@ 8.2.2. Text Input Augmentation
8.2.3. Adaptive Defense.
Backdoor adaptive defense methods that dynamically respond to suspicious behavior have gained traction within the computer vision domain (Guo et al., 2023; Hou et al., 2024; Li et al., 2021b; Chen et al., 2025b). However, such methods are challenging to apply in Natural Language Processing (NLP) due to the discrete nature of language and the high dimensionality of potential trigger patterns. Adaptive defenses for generative NLP tasks remain underexplored (Sun et al., 2023; Zeng et al., 2024), particularly for tasks like Text-to-SQL that involve both the fuzziness of natural language and the structural constraints of SQL outputs. While some studies have proposed dynamic detection mechanisms for SQL injection (Bisht et al., 2010; Bandhakavi et al., 2007), they often introduce significant computational or runtime overhead, and adaptive defenses focused solely on SQL are not applicable to the Text-to-SQL paradigm. For Text-to-SQL tasks, a potential direction (inspired by Zeng et al. (Zeng et al., 2024).), is to analyze the semantic alignment between inputs and outputs in embedding space, and to predefine the semantic boundaries of legitimate SQL queries to dynamically filter out toxic generations.
8.3. Model-level Defense
In addition to filtering and reviewing inputs and outputs, model-level defenses, such as backdoor detection (Liu et al., 2019; Wang et al., 2023) and removal (Wang et al., 2019), are also viable. However, these strategies are typically resource-intensive and costly to implement.
8.3.1. Backdoor Persistence Assessment.
Users often download ready-to-use models from open-source platforms; if these models are already poisoned, can secondary fine-tuning mitigate or eliminate the embedded backdoor? To investigate this, we randomly split the Spider training dataset at a $8:2$ ratio to simulate the attacker’s fine-tuning phase and the user’s subsequent fine-tuning phase. The result, shown in Table 10, reveals that user-led clean fine-tuning does not mitigate the backdoor. In some cases, it can even unintentionally reinforce the attack, further increasing the ASR.
Table 10. Backdoor Persistence Assessment.
Stage Delay-Sudo Comment-double
$S$ EX ASR $S$ EX ASR
Clean 80.84 61.51 - 80.84 61.51 -
Poisoning SFT 79.11 58.51 75.06 79.78 61.80 80.32
Clean Re-SFT 79.88 59.19 77.12 79.38 59.09 77.57
8.3.2. Backdoor Defense.
While many early backdoor defense methods (Wang et al., 2019; Liu et al., 2019, 2018a) are developed for computer vision models, they are often difficult to directly transfer to NLP tasks due to the discrete and context-sensitive nature of language. Beyond input text detection (Qi et al., 2021a; Yang et al., 2021; Gao et al., 2021), existing methods primarily focus on mitigating poisoning without fully retraining the model (Zeng et al., 2024; Min et al., 2024; Sun et al., 2023; Zeng et al., 2022; Liu et al., 2023b). However, such mitigation strategies are often time-consuming and computationally expensive. Moreover, only limited research has addressed the defense of poisoned NLP models involving multiple trigger types (Graf et al., 2024). As a result, there is a pressing need to further investigate defense mechanisms tailored to NLP, particularly for structured and generative tasks like Text-to-SQL. Given these challenges, we recommend prioritizing (1) input and output data filtering, (2) be careful when downloading and deploying Text-to-SQL models, and (3) be careful on dataset for training or fine-tuning, rather than relying on backdoor mitigation as the first line of defense.
@@ 9. CONCLUSION
In this work, we propose ToxicSQL, a framework for implanting configurable backdoors into Text-to-SQL models through poisoned fine-tuning. By predefining malicious payloads–such as SQL injection statements–and lightweight triggers, ToxicSQL enables efficient and stealthy attacks. Remarkably, even single-character triggers can lead to severe data leakage and manipulation risks in database applications. Extensive experiments demonstrate the robustness and transferability of these attacks across different models, payloads, and query styles. These findings highlight the urgent need for the database and natural language processing communities to collaboratively develop adaptive defense mechanisms that go beyond existing filtering methods and provide more fine-grained detection of abnormal behaviors. Future work will explore dynamic and context-aware mitigation and defense strategies to protect Text-to-SQL pipelines from such emerging threats.